OpenStudy (anonymous):

(3x+2)(x-4)=0 .. Use the Zero Product Property ! :)

OpenStudy (chrissy):

-2/3,4

OpenStudy (anonymous):

either 3x + 2 = 0 or x - 4 = 0 solve 3x + 2 = 0 3x = -2 x = -2/3 or x - 4 = 0 x = 4